OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 2.

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

here log has base 16

OpenStudy (anonymous):

32=16^(x+2) now make base =2 both side as 2 is common factor of both

OpenStudy (anonymous):

@Breeziee

OpenStudy (anonymous):

So for 32 I got 2^5

OpenStudy (anonymous):

x =3?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

on right sid e make the base also 2

OpenStudy (anonymous):

16=2^4 2^5=16^(x+2)=2^(4*(x+2)) now comapre the power 5=4*(x+2)

OpenStudy (anonymous):

x = 5/4?

OpenStudy (anonymous):

x+2=5/4

OpenStudy (anonymous):

x=(5/4)-2

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

2^5=16^(x+2)= 2^5=2^(4*(x+2)) comparing power 5=4*(x+2=4x+4*2 5=4x+8 4x=5-8=-3 x=-3/4

OpenStudy (anonymous):

@Breeziee